OpenStudy (anonymous):

What is the slope-intercept form of the equation -3x+2y=12

OpenStudy (anonymous):

do you know about the slope intercept formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright you need to get this equation into slope intercept form y=mx+b how would you do that?

OpenStudy (anonymous):

-3x+2y=12 what would you do first?

OpenStudy (anonymous):

switch 2y with -3x??

OpenStudy (anonymous):

almost you would add -3x on the other side of the equal sign so it shows 2y=3x+12 do you know how to get y by itself?

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

alright how?

OpenStudy (anonymous):

subtract 2 on both sides

OpenStudy (anonymous):

no divide on both sides

OpenStudy (anonymous):

alright lol was gonna ask you about that so now the equation would read y=\[\frac{ 3 }{ 2 }+\frac{ 12 }{ 2 }\] can 12/2 be reduced?

OpenStudy (anonymous):

yes....6

OpenStudy (anonymous):

good so the final equation will look like \[y=\frac{ 3 }{ 2 }x+6\] slope of the line is 3/2 y intercept is 6

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

anytime